# Reinforcement learning to learn quantum states for Heisenberg scaling accuracy

### v1

@@ Abstract
Learning quantum states is a crucial task for realizing the potential of quantum information technology. Recently, neural approaches have emerged as promising methods for learning quantum states. We propose a meta-learning model that employs reinforcement learning (RL) to optimize the process of learning quantum states. For learning quantum states, our scheme trains a Hardware efficient ansatz with a blackbox optimization algorithm, called evolution strategy (ES). To enhance the efficiency of ES, a RL agent dynamically adjusts the hyperparameters of ES. To facilitate the RL training, we introduce an action repetition strategy inspired by curriculum learning. The RL agent significantly improves the sample efficiency of learning random quantum states, and achieves infidelity scaling close to the Heisenberg limit. We showcase that the RL agent trained using $3$-qubit states can be generalized to learning up to $5$-qubit states. These results highlight the utility of RL-driven meta-learning to enhance the efficiency and generalizability of learning quantum states. Our approach can be applicable to improve quantum control, quantum optimization, and quantum machine learning.
@@
Learning quantum states has emerged as a significant problem in the field of quantum computing and machine learning [1, 2]. As a tool for analyzing and benchmarking quantum systems, learning quantum states is essential for realizing quantum information technology [3]. The purpose of learning quantum states is to estimate quantum states with high fidelity based on samples obtained by measuring copies of states. Quantum state tomography (QST) is the standard method for learning quantum states [4, 5], but this method becomes impractical for large systems as the number of measurements required scales exponentially with the size of system. This limitation fuels research directions to devise various strategies for QST using matrix product state [6, 7], permutation invariance of state [8], shadow tomography [9], and Bayesian estimation [10].
@@
With the advent of deep learning, there has been a growing body of work using a neural network-based model for learning quantum states [11, 12, 13, 14, 15, 16, 17, 18, 19]. Generative models provide sample-efficient methods for learning many-body quantum states based on restricted Boltzmann machine [20], variational autoencoder [21], and recurrent neural network [22]. These models are trained with unsupervised learning. Using supervised learning, a model of feed-forward neural network enhances QST by removing errors [23]. The power of these approaches stems from the expressivity of neural network, the ability to represent intricate features of quantum states in terms of a small number of parameters [24]. In the framework of quantum machine learning, a model based on a quantum neural network has also been employed for learning quantum systems [25, 26, 27, 28, 29]. The quantum model can be implemented with parameterized quantum gates [30, 31], making it well suited for learning states of quantum computers. Moreover, it is known that a quantum neural network can have higher expressive power than classical neural networks of comparable size [32, 33].
@@
Meanwhile, the generalizability, the ability to generalize to unexperienced tasks, is one of the hallmarks of machine learning model. However, some of the models for learning quantum states are difficult to generalize as each model is tailored to learning a specific quantum state. Learning models with the generalizability can be found in attempts to improve a quantum optimization algorithm using a deep learning technique [34, 35, 36, 37]. These works employ a hybrid approach that leverages a classical model to learn a quantum model. Specifically, Verdon et al. propose a model based on a recurrent neural network to initialize the parameters of quantum approximate optimization algorithm (QAOA) circuit [35]. Similar work by Wilson et al. uses a recurrent neural network model to learn the gradient of objective function of QAOA [36]. Khairy et al. show that a feed-forward neural network model can be used to learn the hyperparameters of QAOA with reinforcement learning [37]. They show that the hybrid models can be generalized to solving problems of higher dimension than the dimension of problems for training. The deep learning technique in the previous works can be categorized as meta-learning, also called learning to learn, which is employed to enable a model to adapt to unexperienced tasks [38, 39]. Meta-learning methods for learning quantum states have also been proposed, where a classical model is used to learn another classical model [40] or Bayesian estimation algorithm [41]. For the task of learning quantum states, the validity and generalizability of meta-learning models in the hybrid setting remains an area of exploration.
@@
In this work, we propose a meta-learning approach which leverages reinforcement learning (RL) to learn quantum states. The meta-learning model is based on the quantum and classical neural networks. We consider a Hardware efficient ansatz as the quantum neural network model [42]. The quantum model learns a quantum state with a blackbox optimization algorithm, called evolution strategy (ES) [43]. To enhance the process of learning quantum states, a RL agent consisting of classical neural networks dynamically adjusts the hyperparameters of ES. We use single-shot measurements to train the quantum and classical models. It has been shown that using single-shot measurement can exhibit high shot efficiency in learning quantum states [25, 26]. To facilitate the RL, we introduce action repetition strategy inspired by curriculum learning, a technique for training a learning model by gradually increasing difficulty of problems [44]. The trained RL agent significantly improves the shot efficiency of learning quantum states, and enables the infidelity $\bar{f}$ to approach the scaling of Heisenberg limit for the total number of measurement shots ${\cal C}$ [45], i.e., $\bar{f}\sim{\cal O}({\cal C}^{-1})$. We generalize the RL agent to learning quantum states of higher dimension than the dimension of quantum states used for training.
@@
The remainder of this paper is organized as follows. We explain the scheme of meta-learning in Section II. In section III, we present results of meta-learning for random pure states, and introduce the action repetition strategy. We also investigate the generalizability of our model. We give discussion and conclusion in Section IV.
@@
Figure 1: Schematic of reinforcement learning to learn quantum states. (a) In the environment, the evolution strategy trains a Hardware efficient ansatz (HEA) in (b) to learn an input state $\left|\psi\right>$. The observation $o_{t}$ is the success count $\cal C$, and the HEA is trained to increase the success count. The RL agent learns the action $a_{t}$, the hyperparameters of evolution strategy, to enhance the learning quantum state. To enable the learning quantum states to be finished quickly, we give a penalty $r_{t}=-1$ for each time step $t$, and no penalty $r_{t}=0$ for the time step when the learning quantum states is finished. The rollout datasets are generated by a policy-guided tree search, and stored in the data buffer. We use the Actor-Critic algorithm to train the RL agent with a subdataset $\cal B$ randomly drawn from the buffer. (b) HEA circuit. $U_{3}$ is a generic single-qubit gate $\exp({i\vec{\theta}\cdot\vec{\sigma}/2})$, where $\vec{\theta}=(\theta_{x},\theta_{y},\theta_{z})$ and $\vec{\sigma}$ is a vector of Pauli operators $(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})$. For 1-qubit states, we use a single $U_{3}$ gate. For $N$-qubit states, we use a HEA of $L$ layers which has $3N(L+1)$ parameters. (c) Feed-forward neural networks for the RL agent. The input of each network is the observation $o_{t}$. $Q_{\phi_{1}}$ is the observation-value function and $\pi_{\phi_{2}}$ is the policy of action.
@@ II Meta-learning: RL to learn
The schematic of meta-learning model is illustrated in Fig. 1 (a). The environment contains an evolution strategy and a quantum system. As shown in Fig. 1 (b), the quantum system consists of a quantum state $\left|\psi\right>$, a Hardware efficient ansatz (HEA) $\hat{U}(\vec{\theta})$, and a measurement. The evolution strategy trains the HEA to learn an input quantum state based on the measurement. The RL agent learns the hyperparameters of evolution strategy using two classical models in Fig. 1 (c), called actor and critic. We describe our RL scheme with a Markov decision process.
@@
Markov decision process (MDP) is a framework to describe a sequential interaction that takes place between an environment and an agent [46]. The agent takes an action $a$ based on observation $o$ about the environment. As a result of an action, the state of environment is changed, and the agent receives a reward $r$ from the environment. A purpose of agent is to learn the policy of action $\pi$ to maximize cumulative rewards (see Appendix A for a detailed description).
@@
In determining approach of agent for formulating and solving a MDP, observability of the environment is a significant factor to consider. When the elements of MDP are fully known, that is, assuming full observability of environment, an agent can derive the exact policy based on Bellman’s principle of optimality [47]. However, in many scenarios, an agent cannot completely specify the state of environment and its dynamics due to limited (or partial) observability. Partially observable MDP (POMDP) is a generalized framework accommodating such limited observability [48]. To solve a POMDP, an agent can take a heuristic method through an iterative process. A deep RL is known as a method for learning the heuristic policies in the absence of full observability [49].
@@
The problem we consider is the case that the RL agent does not have the full observability of environment. A quantum state transformed by the HEA, $\hat{U}(\vec{\theta})\left|\psi\right>$, represents the state of environment. However, the exact state of environment is unknown to the agent, and the single-shot measurement can extract only a partial information about the quantum state. Thus, we can formulate our problem with the description of POMDP [48, 50, 51] defined by a tuple $\left({\cal S},{\cal A},{\cal T},{\cal R},\Omega,{\cal O}\right)$. The elements of POMDP are explained in Appendix B.
@@
The major challenge in applying RL to quantum systems is the partial observability by quantum measurements [50]. To apply RL for control of quantum systems, there are attempts to utilize additional information about quantum states such as fidelity [52, 53, 54, 55, 56, 57], but computing this quantity is typically expensive, comparable to learning quantum states by the standard QST methods. In contrast, the RL method proposed by Sivak et al. learns the control policy based on observable quantities [51]. In the same spirit, we train our RL agent using the observation $o$ obtained by the measurement. In the following subsections, we describe in detail the configuration of environment, agent, and RL algorithm.
@@ II.1 Environment
Learning a quantum state.—Consider an unknown $N$-qubit state $\left|\psi\right>$ as the input of HEA $\hat{U}(\vec{\theta})$. We measure the quantum state transformed by the HEA, $\hat{U}(\vec{\theta})\left|\psi\right>$, with a measurement $M=\{\hat{M}_{s},\hat{M}_{f}\}$ prepared in known bases, where
@@
The outcome $s$ stands for success and $f$ does fail. We perform the measurement until a fail outcome appears, and count the number of success outcomes before the fail outcome. If a fail outcome appears, we change the parameters of quantum circuit based on the number of success outcomes. Otherwise, we retain the parameters. This measurement scheme is the same as that of single-shot measurement learning (SSML) [25, 26]. While SSML uses a random search algorithm to update the parameters, we use the evolution strategy which will be explained later.
@@
Optimizer.—To train the HEA, we employ an evolution strategy (ES) [58]. An ES consists of following steps: (i) assuming a probability distribution for sampling, (ii) sampling parameters around a current parameter, (iii) evaluating the parameter samples, and (iv) updating the current parameter based on the evaluations. In our problem, the ES uses a multivariate Gaussian distribution to sample parameters [43], and the evaluation corresponds to measuring the success count.
The purpose of ES is to find the parameters of HEA $\vec{\theta}$ to achieve the halting condition ${\cal C}(\vec{\theta})\geq{\cal C}_{\text{target}}$. For this purpose, we consider the expectation of success count as an objective function of ES
@@
1:Set an input quantum state $\left|\psi_{1}\right>$, the HEA $\hat{U}(\vec{\theta}_{0})$, the target success count ${\cal C}_{\text{target}}$, the number of samples $k$, and the RL agent $\pi$
@@
4:     Draw $\sigma_{t}$ and $\eta_{t}$ from the agent, i.e., $(\sigma_{t},\eta_{t})\sim\pi(a_{t}|o_{t})$
5:     Sample $\vec{\epsilon}_{1},\cdots,\vec{\epsilon}_{k}\sim{\cal N}(0,I)$
6:     Measure ${\cal C}\left(\vec{\theta}_{t}+\sigma_{t}\vec{\epsilon}_{i}\right)$ for $i=1,\cdots,k$
7:     if any ${\cal C}\geq{\cal C}_{\text{target}}$ then
8:         break
9:     end if
10:     $\vec{\theta}_{t+1}\leftarrow\vec{\theta}_{t}+\eta_{t}\nabla_{\vec{\theta}}\tilde{J}(\vec{\theta}_{t})$triangleright Eq. (7)
11:     $\left|\psi_{t+1}\right>\leftarrow\hat{U}(\vec{\theta}_{t+1})\left|\psi_{1}\right>$
12:end for
13:${\cal C}_{\text{total}}\leftarrow$ the sum of success counts
14:$t_{H}\leftarrow$ $t$
15:$\vec{\theta}_{\text{train}}\leftarrow\vec{\theta}$ s.t. ${\cal C}(\vec{\theta})\geq{\cal C}_{\text{target}}$
@@
The hyperparameter of sampling range, also called step size, is a significant factor to determine the performance of ES. The sampling range decides the range of search: A larger sampling range induces more exploration on the parameter space. A smaller sampling range focuses on exploitation, searching in a narrower region to enhance solution. In advanced variants of ES such as covariance matrix adaptation ES [59], the sampling range and learning rate are dynamically adjusted, and their combination determines the covariance matrix (or the inverse of Hessian matrix) which reflects the local geometry of objective function. In our scheme, the sampling range $\sigma$ and learning rate $\eta$ are dynamically changed according to the policy $\pi$.
@@
The pseudocode of ES is shown in Algorithm II.1. Unlike the two hyperparameters $\sigma$ and $\eta$, we fix the number of samples $k$ to a single value to keep the statistical uncertainty of the gradient estimator at a certain degree during the execution of ES. ${\cal C}_{\text{total}}$ is the total success count used, and the halting time, $t_{H}$, is the time steps taken until the training is finished. Note that the time here means that how many loops the ES takes until the success count reaches the halting condition ${\cal C}\geq{\cal C}_{\text{target}}$. In a real experiment, the total success count is a quantity which is proportional to physical runtime for the training.
@@
An optimizer typically used to train a quantum neural network is a gradient method [60, 61], but it can suffer the gradient vanishing problem, called barren plateau [62, 63], making the training intractable. While the gradient vanishing effect cannot be fully alleviated [64], recent studies have shown that an ES can be useful for training some cases of quantum neural networks [65, 66]. The gradient methods typically require more evaluation points than parameters to obtain directional derivatives [60]. In contrast, we will show that the ES can train the HEA with the fewer evaluations than the parameters of HEA.
@@
Figure 2: Results of training RL agent for the target success count ${\cal C}_{\text{target}}=10^{4}$. For each RL training episode $T$, we perform learning a thousand random states. $\langle{\cal C}_{\text{total}}\rangle$ is the average of total success count. $\langle\bar{f}\rangle$ is the average infidelity of learned states. $\langle t_{H}\rangle$ is the average halting time steps of evolution strategy. (a) The red, green, and blue solid lines are the training curves for learning $1$-, $2$-, and $3$-qubit states with the $16$ actions. The black solid line is obtained by using the extended action space for learning $2$-qubit states. The dashed lines are baseline values. (b) The policy of trained RL agent.

### v2

@@ Abstract
Learning quantum states is a crucial task for realizing quantum information technology. Recently, neural approaches have emerged as promising methods for learning quantum states. We propose a meta-learning model that utilizes reinforcement learning (RL) to optimize the process of learning quantum states. To improve the data efficiency of the RL, we introduce an action repetition strategy inspired by curriculum learning. The RL agent significantly improves the sample efficiency of learning random quantum states, and achieves infidelity scaling close to the Heisenberg limit. We also show that the RL agent trained using $3$-qubit states can generalize to learning up to $5$-qubit states. These results highlight the utility of RL-driven meta-learning to enhance the efficiency and generalizability of learning quantum states. Our approach can be applied to improve quantum control, quantum optimization, and quantum machine learning.
@@
Learning quantum states has emerged as a significant challenge in the field of quantum computing and machine learning [1, 2]. As a tool for analyzing and benchmarking quantum systems, learning quantum states is essential for advancing quantum information technology [3]. The purpose of learning quantum states is to estimate quantum states with high fidelity based on samples obtained by measurements on multiple copies of a state. Quantum state tomography (QST) is the standard method for the quantum state estimation [4, 5], but this method becomes impractical for large systems since the number of measurements required scales exponentially with the system size. This limitation fuels research directions to devise various strategies for QST using matrix product state [6, 7], permutation invariance of state [8], shadow tomography [9], and Bayesian estimation [10].
@@
With the advent of deep learning, there has been a growing body of work using neural network-based models for learning quantum states [11, 12, 13, 14, 15, 16, 17, 18, 19]. Generative models provide sample-efficient methods for learning many-body quantum states based on restricted Boltzmann machines [20], variational autoencoders [21], and recurrent neural networks [22]. These models are typically trained with unsupervised learning. Using supervised learning, a feed-forward neural network model enhances QST by removing errors [23]. The power of these approaches stems from the expressivity of neural networks, their ability to represent intricate features of quantum states in terms of a small number of parameters [24]. In the framework of quantum machine learning, models based on quantum neural networks have also been employed to learn quantum systems [25, 26, 27, 28, 29]. The quantum models can be implemented with parameterized quantum gates [30, 31], making them well suited for learning states of quantum computers. Moreover, a quantum neural network is known to have higher expressive power than classical neural networks of comparable size [32, 33].
@@
Meanwhile, generalizability, the ability to generalize to unexperienced tasks, is one of the hallmarks of machine learning models. However, some of the models for learning quantum states are difficult to generalize as each model is tailored to learning a specific quantum state. Learning models with generalizability can be found in attempts to improve a quantum optimization algorithms using deep learning techniques [34, 35, 36, 37]. These works employ a hybrid approach that leverages a classical model to learn a quantum model. Specifically, Verdon et al. propose a model based on a recurrent neural network to initialize the parameters of quantum approximate optimization algorithm (QAOA) circuit [35]. Similar work by Wilson et al. use a recurrent neural network model to learn the gradient of objective function in QAOA [36]. Khairy et al. show that a feed-forward neural network model can be used to learn the hyperparameters of QAOA with reinforcement learning [37]. They show that the hybrid models can generalize to solving problems of higher dimension than the dimension of problems for training. The deep learning techniques in the aforementioned works can be categorized as meta-learning, also called learning to learn, which is employed to enable a model to adapt to unexperienced tasks [38, 39]. Meta-learning methods for learning quantum states have also been proposed, where a classical model is used to learn another classical model [40] or Bayesian estimation algorithm [41]. For the task of learning quantum states, the validity and generalizability of meta-learning models in the hybrid setting remain areas of exploration.
@@
In this work, we propose a meta-learning framework that utilizes reinforcement learning (RL) to learn quantum states. Our meta-learning model is based on quantum and classical neural networks. We adopt a hardware efficient ansatz as our quantum neural network model [42]. The quantum model learns quantum states using a blackbox optimization algorithm called evolution strategy (ES) [43]. To enhance the process of learning quantum states, an RL agent consisting of classical neural networks dynamically adjusts the hyperparameters of the ES. To improve the data efficiency of the RL training, we devise action repetition strategy inspired by curriculum learning, a technique for training a learning model by gradually increasing difficulty of problems [44, 45, 46]. Our results demonstrate that the trained RL agent significantly improves the shot efficiency in quantum state learning, enabling the infidelity $\bar{f}$ to scale according to the Heisenberg limit with respect to the total number of measurement shots ${\cal C}$ [47], i.e., $\bar{f}\sim{\cal O}({\cal C}^{-1})$. We also show that our RL agent generalizes to learning quantum states of higher dimensions than those used for training.
@@
The remainder of this paper is organized as follows. We explain the scheme of meta-learning in Section II. In section III, we present results of meta-learning for random pure states, and introduce the action repetition strategy. We also examine the generalizability of our model and compare our results to the standard QST based on maximum likelihood estimation. Finally, we provide discussion and conclusion in Section IV.
@@
Figure 1: Schematic of reinforcement learning to learn quantum states. (a) In the environment, the evolution strategy trains a hardware efficient ansatz (HEA) shown in (b) to learn an input state $\left|\psi\right>$. The observation $o_{t}$ represents the success count $\cal C$, and the HEA is trained to maximize it. The RL agent learns the action $a_{t}$, the hyperparameters of the evolution strategy. To expedite the training process, we give a penalty $r_{t}=-1$ to the RL agent at each time step $t$, and no penalty $r_{t}=0$ for the final time step. The rollout datasets are generated using a policy-guided tree search and stored in the data buffer. We use the Actor-Critic algorithm to train the RL agent with a subdataset $\cal B$ randomly drawn from the buffer. (b) HEA circuit. $U_{3}$ is a generic single-qubit gate $\exp({i\vec{\theta}\cdot\vec{\sigma}/2})$, where $\vec{\theta}=(\theta_{x},\theta_{y},\theta_{z})$ and $\vec{\sigma}$ denotes the vector of Pauli operators $(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})$. For 1-qubit states, we use a single $U_{3}$ gate. For $N$-qubit states, we use the HEA of $L$ layers, consisting of $3N(L+1)$ trainable parameters. (c) Feed-forward neural networks for the RL agent. Each network adopts the observation $o_{t}$ as input. $Q_{\phi_{1}}$ is the observation-value function and $\pi_{\phi_{2}}$ is the policy of action.
@@ II Meta-learning: RL to learn
A schematic of the meta-learning model is illustrated in Fig. 1 (a). The environment consists of an evolution strategy and a quantum system. As depicted in Fig. 1 (b), the quantum system comprises a quantum state $\left|\psi\right>$, a hardware efficient ansatz (HEA) $\hat{U}(\vec{\theta})$, and a measurement. The evolution strategy optimizes the HEA to learn an input quantum state based on measurement outcomes. The RL agent learns the hyperparameters of the evolution strategy using two classical models in Fig. 1 (c), the actor and critic.
@@
We formulate our RL scheme using a Markov decision process. The Markov decision process (MDP) is a framework to describe a sequential interaction between an agent and its environment [48]. The agent takes an action $a$ based on observation $o$ of the environment. The action alters the state of environment, and the agent receives a reward $r$ in response. The goal of the agent is to learn a policy $\pi$ that maximizes cumulative rewards (see Appendix A for details).
@@
In determining approach of agent to formulating and solving a MDP, observability of the environment is a significant factor to consider. When the elements of an MDP are fully known, that is, assuming full observability of environment, an agent can derive the exact policy based on Bellman’s principle of optimality [49]. However, in many scenarios, an agent cannot completely specify the state of the environment or its dynamics due to limited (or partial) observability. A partially observable MDP (POMDP) extends the standard MDP framework to account for such limited observability [50]. To solve the POMDP, an agent can take a heuristic method through an iterative process. Deep RL is widely used to learn heuristic policies under the limited observability [51].
@@
The problem we consider is the case that the RL agent lacks full observability of the environment. The quantum state transformed by the HEA, $\hat{U}(\vec{\theta})\left|\psi\right>$, represents the state of the environment. However, the agent does not have full knowledge of the exact state of the environment, and quantum measurements can only extract partial information about the quantum state. Thus, we can formulate our problem as a POMDP [50, 52, 53], defined by a tuple $\left({\cal S},{\cal A},{\cal T},{\cal R},\Omega,{\cal O}\right)$. The elements of the POMDP are explained in Appendix B.
@@
A major challenge in applying RL to quantum systems is the partial observability due to quantum measurements [52]. In applying RL to control of quantum systems, there are attempts to utilize additional information about quantum states such as fidelity [54, 55, 56, 57, 58, 59], but computing this quantity is typically expensive, comparable to learning quantum states by the standard QST methods. In contrast, the RL method proposed by Sivak et al. learns a control policy based on observable quantities [53]. In the same spirit, we train our RL agent using the observation $o$ obtained from the measurement. In the following subsections, we provide a detailed description of the environment, agent, and RL algorithm.
@@ II.1 Environment
Learning a quantum state.—Consider an unknown $N$-qubit state $\left|\psi\right>$ as an input of the HEA $\hat{U}(\vec{\theta})$. We measure the quantum state transformed by the HEA, $\hat{U}(\vec{\theta})\left|\psi\right>$, with a measurement $M=\{\hat{M}_{s},\hat{M}_{f}\}$ prepared in known bases, where
@@
The outcome $s$ stands for success and $f$ does fail. We perform the measurement until a fail outcome appears, and count the number of consecutive success outcomes before the fail outcome. If a fail outcome appears, we change the parameters of quantum circuit based on the number of success outcomes. Otherwise, we retain the parameters. This measurement scheme is the same as that of single-shot measurement learning (SSML) [25, 26]. While the SSML uses a weighted random search algorithm to update the parameters, we use the evolution strategy which will be explained later.
@@
Optimizer.—To train the HEA, we employ an evolution strategy (ES) [60]. An ES consists of following steps: (i) assuming a probability distribution for sampling, (ii) sampling parameters around a current parameter, (iii) evaluating the parameter samples, and (iv) updating the current parameter based on the evaluations. In our problem, the ES uses a multivariate Gaussian distribution to sample parameters [43], and the evaluation corresponds to the measurement of success count.
The purpose of ES is to find the parameters of HEA $\vec{\theta}$ to achieve the halting condition ${\cal C}(\vec{\theta})\geq{\cal C}_{\text{target}}$. For this purpose, we consider the expectation of success count as an objective function of our ES
@@
1:Set an input quantum state $\left|\psi_{1}\right>$, the HEA $\hat{U}(\vec{\theta}_{0})$, the target success count ${\cal C}_{\text{target}}$, the action repetition time $t_{\text{rep}}$, the number of samples $k$, and the RL agent $\pi$
@@
4:     if $t\%t_{\text{rep}}=0$ then
5:         Draw $\sigma_{t}$ and $\eta_{t}$ from the agent, $(\sigma_{t},\eta_{t})\sim\pi(a_{t}|o_{t})$
6:     end if
7:     Sample $\vec{\epsilon}_{1},\cdots,\vec{\epsilon}_{k}\sim{\cal N}(0,I)$
8:     Measure ${\cal C}\left(\vec{\theta}_{t}+\sigma_{t}\vec{\epsilon}_{i}\right)$ for $i=1,\cdots,k$
9:     if any ${\cal C}\geq{\cal C}_{\text{target}}$ then
10:         break
11:     end if
12:     $\vec{\theta}_{t+1}\leftarrow\vec{\theta}_{t}+\eta_{t}\nabla_{\vec{\theta}}\tilde{J}(\vec{\theta}_{t})$triangleright Eq. (7)
13:     $\left|\psi_{t+1}\right>\leftarrow\hat{U}(\vec{\theta}_{t+1})\left|\psi_{1}\right>$
14:end for
15:${\cal C}_{\text{total}}\leftarrow$ the sum of success counts
16:$t_{H}\leftarrow$ $t$
17:$\vec{\theta}_{\text{train}}\leftarrow\vec{\theta}$ s.t. ${\cal C}(\vec{\theta})\geq{\cal C}_{\text{target}}$
@@
The sampling range, also called step size, is a key hyperparameter that significantly impacts the performance of ES. The sampling range decides the range of search: A larger sampling range induces more exploration on the parameter space. A smaller sampling range focuses on exploitation, searching in a narrower region to enhance solution. In advanced ES variants, such as covariance matrix adaptation ES [61], both the sampling range and learning rate are dynamically adjusted. Their interplay determines the covariance matrix, which captures the local geometry of objective function. In our scheme, the sampling range $\sigma$ and learning rate $\eta$ are dynamically changed according to the policy $\pi$.
@@
The pseudocode of ES is shown in Algorithm II.1. Unlike the two hyperparameters $\sigma$ and $\eta$, we fix the number of samples $k$ to a single value to keep the statistical uncertainty of the gradient estimator at a certain degree throughout the execution of ES. Each action drawn by the agent is applied for $t_{\text{rep}}$ consecutive time steps, referred to as the action repetition time. The effectiveness of the action repetition will be discussed in the next section. ${\cal C}_{\text{total}}$ is the total success count used, and the halting time, $t_{H}$, is the time steps taken until the training is finished. Note that the time here refers to the number of ES iterations required for the success count to reach the halting condition ${\cal C}\geq{\cal C}_{\text{target}}$. In a real experiment, the total success count is a quantity which is proportional to the physical runtime of the training.
@@
An optimizer typically used to train a quantum neural network is a gradient method [62, 63], but it can suffer the gradient vanishing problem, called barren plateau [64, 65], making the training intractable. While the gradient vanishing effect cannot be fully alleviated [66], recent studies have shown that ESs can be useful for training some cases of quantum neural networks [67, 68]. The gradient methods typically require more evaluation points than parameters to obtain directional derivatives [62]. We will show that the ES can train the HEA with the fewer evaluation points than the parameters of HEA.
@@
Figure 2: Results of training RL agent for the target success count ${\cal C}_{\text{target}}=10^{4}$. For each RL training episode $T$, we perform learning of a thousand random states. $\langle{\cal C}_{\text{total}}\rangle$ is the average of total success count. $\langle\bar{f}\rangle$ is the average infidelity of learned states. $\langle t_{H}\rangle$ is the average halting time steps of evolution strategy. (a) The red, green, and blue solid lines are the training curves for learning $1$-, $2$-, and $3$-qubit states with the $16$ actions. The black solid line is obtained by using the extended action space for learning $2$-qubit states. The dashed lines are baseline values. (b) The policy of trained RL agent.